An applicant for a flight steward stands five feet and two inches tall. Can this applicant qualify if the height qualification is 165 centimeters?


1
Expert's answer
2021-09-07T23:27:57-0400

1 foot = 30.48 centimetres

= 30.48 × 5 = 152.4 cm

I inch = 2.54 centimetres

= 2.54 × 2= 5.08 cm

= 152.4 + 5.08 = 157.1cm

Therefore, the applicant will not qualify